OpenStudy (anonymous):

Once you have found one zero, use Quadratic Formula to find the other two zeros. the first zero is 2 the equation is X^3 + 6x^2+8x - 48

OpenStudy (anonymous):

$$\begin{align*}x^3+6x^2+8x-48&=x^2(x+6)+8(x-6)\\&=x^2(x-2+8)+8(x-2-4)\\&=x^2(x-2)+8(x-2)+8x^2-32\\&=(x-2)(x^2+8)+8(x^2-4)\\&=(x-2)(x^2+8)+4(x-2)(x+2)\\&=(x-2)(x^2+8+4(x+2))\\&=(x-2)(x^2+8+4x+8)\\&=(x-2)(x^2+4x+16)\end{align*}$$Divide through by \(x-2\), which is the factor that made \(x=2\) a zero:$$\frac{x^3+6x^2+8x-48}{x-2}=x^2+4x+16$$Now, we use the quadratic formula:$$\begin{align*}x&=\frac{-4\pm\sqrt{16-4(16)}}2\\&=-2\pm\sqrt{1-4}\\&=-2\pm3i\end{align*}$$

OpenStudy (anonymous):

Ooops, I divided incorrectly... hmm